# Unifying Unsupervised Graph-Level Anomaly Detection and Out-of-Distribution Detection: A Benchmark

### v1

@@ 4.1 Performance Comparison (RQ1)
Experiment Design. We conduct a comprehensive comparison of the detection performance in terms of AUROC, AUPR, and FPR95 of 16 benchmark algorithms on 35 benchmark datasets. For each method and dataset, we conduct 5 runs of experiments and report the average performance.
Experimental Results. Table 3 shows the performance comparison in terms of AUROC, and Fig. 2 provides box plots to overview the model performance across 35 datasets for all metrics. In Appendix E, detailed performance with standard deviation is demonstrated in Table 8, and the box plots of the ranking of each method are demonstrated in Fig. 6. We have the following observations.
Observation ❶: The SOTA GLAD/GLOD methods show excellent performance on both tasks. The results in Table 3 highlight the excellent performance of the SOTA GLAD/GLOD methods on both detection tasks. Specifically, the GLAD methods, SIGNET and OCGTL, demonstrate outstanding performance in OOD detection tasks, achieving competitive results in several datasets. Meanwhile, the GLOD method GOOD-D, while not attaining the best performance in anomaly detection tasks, still performs commendably with an average ranking of 4.60, placing it among the top performers. This observation highlights the intercommunity between GLAD and GLOD tasks, emphasizing the need to apply powerful methods designed for one task to the other.
Observation ❷: No universally superior method. Table 3 demonstrates that no single method consistently outperforms others on more than 12 (out of 35) datasets. Even the top-ranked method, SIGNET, can exhibit sub-optimal performance on several datasets belonging to different types, such as HSE, DD, ES→MU, and ZINC-Size. In Fig. 2, although SIGNET has a higher median and a compact interquartile range, the top 25% of its performance is still lower than that of some other models. Similar unstable performance can also be found in other competitive methods, such as OCGTL, CVTGAD, and GOOD-D. This finding illustrates the diversity of our benchmark datasets, underscoring the challenge of identifying a “universal method” that performs well across all datasets.
@@
Observation ❹: End-to-end methods show consistent superiority over 2-step methods. While 2-step methods show notable performance in certain scenarios, end-to-end methods consistently outperform them. Specifically, in Table 3, the average rankings of most end-to-end methods are below 8, while all 2-step methods have rankings above 8. This consistent superiority highlights the advantages of integrated and unified learning approaches over segmented and two-step processes.
4.2 Generalizability on Near-OOD and Far-OOD (RQ2)
Experiment Design. To evaluate the generalizability of GLAD/GLOD methods in handling near and far OOD samples, we consider two different settings to define near-OOD and far-OOD: (A) intra-inter dataset setting and (B) size-based distance setting. In setting A, we define the intra-dataset samples with different class labels as the near-OOD, while samples from another dataset are considered far-OOD. In setting B, the size of graphs serves as the measure to divide near and far OOD. We redefine the size of training/testing sets for fair comparison. Please refer to Appendix C.3 for a more detailed description of experimental settings.
@@ 4.2 Generalizability on Near-OOD and Far-OOD (RQ2)
Observation ❺: Near-OOD samples are harder to detect compared to far-OOD samples. From Fig. 3, it is evident that end-to-end models generally perform better in detecting far-OOD samples than near-OOD samples. This trend is consistent across various datasets.
Observation ❻: Poor generalizability of several GLAD/GLOD methods in specialized scenarios. Despite the overall effectiveness of GLAD/GLOD methods, their performance has notable limitations, particularly in specialized scenarios. Firstly, in setting A, GLOD approaches (i.e., GOOD-D and GraphDE) exhibit significant performance gaps between near-OOD and far-OOD conditions. This suggests that when OOD samples are similar to ID samples, the detection capability of GLOD methods is significantly compromised. This limitation is even more pronounced in setting B, where GLAD methods significantly underperform 2-step methods.
@@ 4.3 Robustness under Training Set Perturbation (RQ3)
Experiment Design. In this study, we investigate the robustness of GLAD/GLOD methods against the perturbation of the normal/ID training set by contaminating anomaly/OOD samples. Specifically, we set the perturbation ratios as 0%, 10%, 20%, and 30% to explore the impacts of different perturbation strengths. For more details, please refer to Appendix C.3.
@@
Observation ❼: Performance degradation with increasing contamination ratio. From Fig. 4, it is obvious that the performance of GLAD/GLOD methods generally deteriorates as the proportion of OOD samples in the ID training set increases. This trend is evident across multiple datasets and methods. This decline in performance suggests that the presence of anomaly/OOD samples in the training data introduces noise and confounds the models, making it increasingly difficult for them to distinguish between generalized ID and OOD samples during testing.
@@ 4.4 Efficiency Analysis (RQ4)
Experiment Design. We evaluate the computational efficiency of GLAD/GLOD methods using default hyperparameter settings. Our assessment focuses on two main aspects: time efficiency and memory usage. See Appendix C.2 for a detailed description.
@@
Observation ❾: Certain end-to-end methods outperform 2-step methods in terms of both performance and computational costs. Observation ❹ demonstrates that end-to-end methods usually outperform the 2-step methods. In addition to their superior performance, most end-to-end methods exhibit comparable time efficiency and significantly better memory efficiency. As shown in Fig. 5, end-to-end methods achieve optimal results much faster than 2-step methods, except for GOOD-D and GraphDE. In contrast, the iF detector in 2-step methods significantly increases the time required to reach optimal performance, leading to substantially higher time costs compared to most existing methods. In terms of memory usage, as illustrated in Fig. 5, end-to-end methods show much lower CPU and GPU consumption than 2-step methods. Although graph kernel methods do not utilize GPU resources, their high CPU consumption is a critical consideration. To sum up, the majority of end-to-end methods may be preferable for GLAD/GLOD tasks due to their advantages in effectiveness and efficiency.
@@ 5 Conclusion and Future Directions
In this paper, we introduce a unified benchmark, UB-GOLD, for graph-level anomaly detection (GLAD) and graph-level out-of-distribution detection (GLOD), which comprehensively compare the performance of 16 GLAD/GLOD methods on 35 datasets across various scenarios and domains. Based on UB-GOLD, we conduct extensive experiments to analyze the effectiveness, generalizability, robustness, and efficiency of these methods. Observation ❶ highlights the intercommunity between GLAD and GLOD tasks. It reveals that many approaches show comparable performance across both tasks, suggesting a significant overlap in the underlying principles and techniques used for detecting anomalies and OOD samples. Several insightful observations are summarized from the results, providing an in-depth understanding of existing methods and inspiration for future work. Our experiments reveal that SOTA methods for GLAD and GLOD exhibit superior performance across both tasks, underscoring the strong interconnection between these two forefront research domains.
@@
Limitation. Since UB-GOLD mainly focuses on unsupervised GLAD/GLOD tasks, the methods that require class labels (e.g. AAGOD [14]), anomaly labels (e.g. i-GAD [4]), and domain-specific pre-trained models (e.g., PGR-MOOD [11]) are not included for fair comparison. As a long-term evolving project, UB-GOLD will include these methods in the codebase for quick implementation.
Checklist
1.
For all authors…
Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes]
Did you describe the limitations of your work? [Yes] See Section 5: Conclusion and Future Directions
Did you discuss any potential negative societal impacts of your work? [N/A]
Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes]
2.
If you are including theoretical results…
Did you state the full set of assumptions of all theoretical results? [N/A]
Did you include complete proofs of all theoretical results? [N/A]
3.
If you ran experiments (e.g. for benchmarks)…
Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes] See Abstract.
Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes] See Section 3.3:Evaluation Metrics and Implementation Details.
Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [Yes] See Appendix D:Additional Experimental Results.
Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] See Appendix C.2: Additional Experimental Detail.
4.
If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
If your work uses existing assets, did you cite the creators? [Yes]
Did you mention the license of the assets? [Yes] See Appendix D: Reproducibility
Did you include any new assets either in the supplemental material or as a URL? [Yes] See Appendix D: Reproducibility
Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [Yes] See Appendix D: Reproducibility, and cite their paper.
Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [Yes] See Appendix D: Reproducibility
5.
If you used crowdsourcing or conducted research with human subjects…
Did you include the full text of instructions given to participants and screenshots, if applicable? [N/A]
Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A]
Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [N/A]
Appendix A Detailed Description of datasets in UB-GOLD
The description of datasets in UB-GOLD is given as follows.
⚫ Type I: Datasets with intrinsic anomaly.111 Tox21 Challenge Data.
HSE [29]: Focuses on human stress response elements, essential for studying stress-induced cellular changes.
MMP [29]: Targets matrix metalloproteinases, crucial for understanding processes like cancer metastasis.
p53 [29]: Involves the p53 protein, a key player in cancer prevention and apoptosis.
PPAR-gamma [29]: Examines peroxisome proliferator-activated receptor gamma, important for metabolic and inflammatory processes.
⚫ Type II: Datasets with class-based anomaly.222 TUDataset.
@@ Appendix A Detailed Description of datasets in UB-GOLD
COLLAB [35]: Scientific collaboration networks, representing authors and their co-authored papers in various scientific fields.
@@
IMDB-BINARY [35]: Movie collaboration networks, where nodes represent actors and edges denote their co-appearance in films. This dataset helps in analyzing the collaboration patterns within the film industry.
@@
REDDIT-BINARY [35]: Discussion threads from Reddit, each graph representing a thread with nodes as users and edges as interactions. It captures the structure of online discussions and their dynamics.
@@
ENZYMES [36]: Protein tertiary structures, each graph represents an enzyme with nodes as secondary structure elements and edges as spatial adjacencies, crucial for biochemical and functional studies.
@@
PROTEINS [37]: Comprehensive protein structures and interaction data, where nodes represent amino acids and edges denote interactions, providing insights into protein functions and interactions.
@@
DD [20]: Protein-protein interaction networks, capturing the intricate relationships between proteins within biological systems, essential for understanding cellular functions.
@@
BZR [38]: Benzodiazepine receptor ligands, with nodes representing atoms and edges representing bonds, used for studying molecular interactions with benzodiazepine receptors.
@@
AIDS [30]: Chemical compounds screened for antiviral activity against HIV, where each graph represents a molecule, useful in drug discovery and antiviral research.
@@
COX2 [30] : Cyclooxygenase-2 inhibitors, where nodes and edges represent molecular structures, important for studying anti-inflammatory drug properties.
@@
NCI1 [39]: Chemical compounds screened for anti-cancer activity, each graph representing a molecule, used for evaluating potential anti-cancer drugs.
@@
DHFR [30]: Dihydrofolate reductase inhibitors, with graphs representing molecular structures, focusing on compounds inhibiting the DHFR enzyme, vital for cancer and bacterial infection treatments.
@@
All graphs in the following dataset represent molecules, where nodes are atoms, and edges are chemical bonds.
BBBP [40]: BBBP is the Blood–brain barrier penetration (BBBP) dataset includes binary labels for over 2000 compounds on their permeability properties.
BACE [41]: BACE provides a series of human -secretases as well as their binary label, and all data are experimental values reported in the scientific literature over the last decade.
CLINTOX [42]: The ClinTox dataset compares drugs that have been approved by the FDA with drugs that have failed in clinical trials for toxicity reasons.
LIPO [38]: The full name of LIPO is Lipophilicity, which is an important feature of drug molecules that affects both membrane permeability and solubility.
FREESOLV [43]: The Free Solvation database (FreeSolv) provides experimental and calculated free energies of hydration of small molecules in water.
TOXCAST [44]: ToxCast providing toxicology data for a library of compounds based on high-throughput screening and includes qualitative results of over 600 experiments on 8615 compounds.
ESOL [45]: ESOL is a small dataset containing water solubility data for 1128 compounds with the goal of estimating solubility from chemical structures.
MUV [46]: The Maximum Unbiased Validation (MUV) group is another benchmark dataset selected from PubChem BioAssay by applying a modified nearest neighbor analysis
TOX21 [38]: The dataset contains qualitative toxicity measurements of 8014 compounds against 12 different targets, including nuclear receptors and stress response pathways.
SIDER [47]: The Side Effect Resource (SIDER) is a database of marketed drugs and adverse drug reactions which measured for 1427 approved drugs.
PTC-MR [48]: PTC dataset labels compounds based on their carcinogenicity where MR Indicates that their rodent is a male rat.
@@
GOOD [19]: A systematic benchmark specifically tailored for the graph OOD problem. We utilize two molecular datasets for OOD detection tasks: (1) GOOD-HIV is a small-scale real-world molecular dataset which aim to predict whether this molecule can inhibit HIV replication. (2) GOOD-ZINC is a real-world molecular property regression dataset from ZINC database and aims at predicting molecular solubility. Each dataset comprises two ID-OOD splitting strategies (scaffold and size), resulting in a total of 4 distinct datasets.
DrugOOD [12]: This OOD benchmark is designed for AI-aided drug discovery. It includes three ID-OOD splitting strategies: assay, scaffold, and size. These strategies are applied to two measurements (IC50 and EC50), resulting in six datasets. Each dataset comprises a binary classification task aimed at predicting drug target binding affinity.
Appendix B Detailed Description of Algorithms in UB-GOLD

### v2

@@ 4.1 Performance Comparison (RQ1)
Experiment design. We conduct a comprehensive comparison of the detection performance in terms of AUROC, AUPR, and FPR95 of 18 benchmark algorithms on 35 benchmark datasets. For each method and dataset, we conduct 5 runs of experiments and report the average performance.
Experimental results. Table 3 shows the performance comparison in terms of AUROC, and Fig. 3 provides box plots to overview the model performance across 35 datasets for all metrics. In Appendix F, detailed performance with standard deviation is demonstrated in Table 8, and the box plots of the ranking of each method are demonstrated in Fig. 7. We have the following observations.
Observation ❶: The SOTA GLAD/GLOD methods show excellent performance on both tasks. The results in Table 3 highlight the excellent performance of the SOTA GLAD/GLOD methods on both detection tasks. Specifically, the GLAD methods, SIGNET, CVTGAD, and OCGTL, demonstrate outstanding performance in OOD detection tasks, achieving competitive results in several datasets. Meanwhile, the GLOD method GOOD-D, while not attaining the best performance in anomaly detection tasks, still performs commendably with an average ranking of 4.14, placing it among the top performers. This observation highlights the intercommunity between GLAD and GLOD tasks, emphasizing the need to apply powerful methods designed for one task to the other.
Observation ❷: No universally superior method. Table 3 demonstrates that no single method consistently outperforms others on more than 12 (out of 35) datasets. Even the top-ranked method, SIGNET, can exhibit sub-optimal performance on several datasets belonging to different types, such as HSE, DD, ES→MU, and ZINC-Size. In Fig. 3, although SIGNET has a higher median and a compact interquartile range, the top 25% of its performance is still lower than that of some other models. Similar unstable performance can also be found in other competitive methods, such as OCGTL, CVTGAD, and GOOD-D. This finding illustrates the diversity of our benchmark datasets, underscoring the challenge of identifying a “universal method” that performs well across all datasets. This is particularly evident in the AUPRC metric. For instance, on HSE, SIGNET achieves an AUROC of 64.56 vs. 71.42 (SOTA). On DD, it scores 74.53 vs. 80.76 (SOTA). In the inter-dataset shift ES→MU, it achieves 91.43 vs. 99.64 (SOTA), and in the intra-dataset shift ZINC-Size, it scores 57.29 vs. 68.43 (SOTA).
@@
Observation ❹: End-to-end methods show consistent superiority over two-step methods. While two-step methods show notable performance in certain scenarios, end-to-end methods consistently outperform them. Specifically, in Table 3, the average rankings of most end-to-end methods are below 10, while all two-step methods have rankings above 10. This consistent superiority highlights the advantages of integrated and unified learning approaches over segmented and two-step processes.
4.2 OOD sensitivity spectrum on Near-OOD and Far-OOD (RQ2)
Experiment design. To evaluate the OOD sensitivity spectrum of GLAD/GLOD methods in handling near and far OOD samples, we consider two different settings to define near-OOD and far-OOD: (A) intra-inter dataset setting and (B) size-based distance setting. In setting A, we define the intra-dataset samples with different class labels as the near-OOD, while samples from another dataset are considered far-OOD. In setting B, the size of graphs serves as the measure to divide near and far OOD. We redefine the size of training/testing sets for fair comparison. Please refer to Appendix D.4 for a more detailed description of experimental settings.
@@ 4.2 OOD sensitivity spectrum on Near-OOD and Far-OOD (RQ2)
Observation ❺: Near-OOD samples are harder to detect compared to far-OOD samples. From Fig. 4, it is evident that end-to-end models generally perform better in detecting far-OOD samples than near-OOD samples. This trend is consistent across various datasets. For instance, in the AIDS dataset, models like GOOD-D, GraphDE, and GOODAT achieve highly better AUROC values for far-OOD conditions than for near-OOD. Similar patterns are observed in the other datasets, where most models display better performance in far-OOD scenarios. This consistent performance discrepancy underscores the challenge of detecting near-OOD samples that closely resemble the ID data, highlighting the need for enhanced model sensitivity to subtle deviations.
Observation ❻: Poor OOD sensitivity spectrum of several GLAD/GLOD methods in specialized scenarios. Despite the overall effectiveness of GLAD/GLOD methods, their performance has notable limitations, particularly in specialized scenarios. Firstly, in setting A, GLOD approaches (i.e., GOOD-D and GraphDE) exhibit significant performance gaps between near-OOD and far-OOD conditions. This suggests that when OOD samples are similar to ID samples, the detection capability of GLOD methods is significantly compromised. This limitation is even more pronounced in setting B, where GLAD methods significantly underperform two-step methods. In these scenarios, GLAD methods struggle with size-based deviations, resulting in substantial performance gaps.
@@ 4.3 Robustness under Training Set Perturbation (RQ3)
Experiment design. In this study, we investigate the robustness of GLAD/GLOD methods against the perturbation of the normal/ID training set by contaminating anomaly/OOD samples. Specifically, we set the perturbation ratios as 0%, 10%, 20%, and 30% to explore the impacts of different perturbation strengths. For more details, please refer to Appendix D.4.
@@
Observation ❼: Performance degradation with increasing contamination ratio. From Fig. 5, it is obvious that the performance of GLAD/GLOD methods generally deteriorates as the proportion of OOD samples in the ID training set increases. This trend is evident across multiple datasets and methods. For example, in the PROTEINS, BZR, and COX2 datasets, we observe that models such as GLocalKD, OCGTL, SIGNET, GOOD-D, and others show a consistent decrease in AUROC values as the proportion of OOD contamination increases from 0% to 30%. This decline in performance suggests that the presence of anomaly/OOD samples in the training data introduces noise and confounds the models, making it increasingly difficult for them to distinguish between generalized ID and OOD samples during testing.
@@ 4.4 Efficiency Analysis (RQ4)
Experiment design. We evaluate the computational efficiency of GLAD/GLOD methods using default hyperparameter settings. Our assessment focuses on two main aspects: time efficiency and memory usage. See Appendix D.3 for a detailed description.
@@
Observation ❾: Certain end-to-end methods outperform two-step methods in terms of both performance and computational costs. Observation ❹ demonstrates that end-to-end methods usually outperform the two-step methods. In addition to their superior performance, most end-to-end methods exhibit comparable time efficiency and significantly better memory efficiency. As shown in Fig. 6, end-to-end methods achieve optimal results much faster than two-step methods, except for GOOD-D and GraphDE. In contrast, the iF detector in two-step methods significantly increases the time required to reach optimal performance, leading to substantially higher time costs compared to most existing methods. In terms of memory usage, as illustrated in Fig. 6, end-to-end methods show much lower CPU and GPU consumption than two-step methods. Although graph kernel methods do not utilize GPU resources, their high CPU consumption is a critical consideration. To sum up, the majority of end-to-end methods may be preferable for GLAD/GLOD tasks due to their advantages in effectiveness and efficiency.
@@ 5 Conclusion and Future Directions
In this paper, we introduce a unified benchmark, UB-GOLD, for graph-level anomaly detection (GLAD) and graph-level out-of-distribution detection (GLOD), which comprehensively compare the performance of 18 GLAD/GLOD methods on 35 datasets across various scenarios and domains. Based on UB-GOLD, we conduct extensive experiments to analyze the effectiveness, OOD sensitivity spectrum, robustness, and efficiency of these methods. Additionally, we provide a visualization and analysis of the Judge Score Distribution in the Appendix F.3. Observation ❶ highlights the intercommunity between GLAD and GLOD tasks. It reveals that many approaches show comparable performance across both tasks, suggesting a significant overlap in the underlying principles and techniques used for detecting anomalies and OOD samples. Several insightful observations are summarized from the results, providing an in-depth understanding of existing methods and inspiration for future work. Our experiments reveal that SOTA methods for GLAD and GLOD exhibit superior performance across both tasks, underscoring the strong interconnection between these two forefront research domains.
@@
Limitation. Since UB-GOLD mainly focuses on unsupervised GLAD/GLOD tasks, the methods that require anomaly labels (e.g. i-GAD (Zhang et al., 2022)), and domain-specific pre-trained models (e.g., PGR-MOOD (Shen et al., 2024)) are not included for a fair comparison. As a long-term evolving project, UB-GOLD will include these methods in the codebase for quick implementation.
Ethics Statement
This research is dedicated solely to scientific inquiry without involving human subjects, animals, or materials that may pose environmental concerns. As such, we do not anticipate any ethical risks or conflicts of interest. We are committed to upholding the highest standards of scientific integrity and ethics to ensure the accuracy and credibility of our findings.
ACKNOWLEDGMENTS
This work was supported by a grant from the National Natural Science Foundation of China under grants (No.62372211, 62272191), the Foundation of the National Key Research and Development of China (No.2021ZD0112500), the International Science and Technology Cooperation Program of Jilin Province (No.20230402076GH, No. 20240402067GH), and the Science and Technology Development Program of Jilin Province (No. 20220201153GX). Y. Liu and S. Pan were partially supported by Australian Research Council (ARC) under grant DP240101547.
Appendix A Why UB-GOLD focuses on unsupervised scenarios
Node-level anomaly detection (AD) and out-of-distribution (OOD) detection can be performed in both supervised (Tang et al., 2023) and unsupervised (Liu et al., 2022) settings due to their transductive nature. In node-level tasks, the goal is to detect anomalous or OOD nodes within a single graph. Since the entire graph is observed, labeled data is often accessible, and supervised methods can be effectively applied. Supervised approaches leverage these labels to guide the model in identifying anomalies, while unsupervised methods detect deviations from normal node behavior without labels. Both approaches can coexist, as node-level tasks benefit from the complete visibility of the graph during training.
In contrast, graph-level AD/OOD detection poses different challenges, making supervised or semi-supervised approaches less meaningful. Unlike node-level tasks, graph-level detection deals with unseen, independent graphs at test time, where labeled anomaly or OOD samples are rarely available in real-world scenarios. Relying on labeled data in supervised or semi-supervised settings shifts the focus to classification rather than detecting unexpected anomalies. This undermines the core challenge of generalizing across unseen graphs. Furthermore, graph-level tasks share a closer resemblance to computer vision (CV) problems, where the primary challenge is detecting shifts in distribution without labeled anomalies. Our inspiration stems from the success of benchmarks like OpenOOD (Zhang et al., 2023) in CV, where unsupervised methods have proven to be more effective at handling distribution shifts in the absence of labeled OOD data.
By focusing solely on unsupervised methods in UB-GOLD, we ensure that the benchmark reflects the true nature of graph-level AD/OOD detection in real-world applications. Unsupervised approaches eliminate the dependency on labeled data, allowing models to learn from the normal distribution and detect deviations more naturally. This aligns with practical conditions, where labeled OOD or anomaly samples are rarely available. Additionally, an unsupervised framework promotes more robust generalization across diverse datasets, pushing the boundaries of AD/OOD detection in the graph domain, similar to the advancements driven by OpenOOD in computer vision.
Appendix B Detailed Description of datasets in UB-GOLD
The description of datasets in UB-GOLD is given as follows. To ensure the reliability of our benchmark and avoid any data leakage, we carefully curated all dataset splits to guarantee no overlap between ID and OOD samples, or between training and test sets. Each dataset was meticulously prepared to maintain a clear separation, ensuring that OOD samples are entirely distinct from ID data. This strict partitioning is consistently applied across all experimental setups, providing a robust and unbiased evaluation of the methods.
⚫ Type I: Datasets with intrinsic anomaly.222 Tox21 Challenge Data.
Datasets under this type include real anomalies inherently present in the data, defined based on the biological or chemical properties of the samples. These anomalies are characterized by rare or atypical patterns that deviate from the majority of the data. Such intrinsic anomalies are valuable for evaluating the model’s ability to detect deviations from the norm, a common challenge in real-world data analysis.
HSE (Abdelaziz et al., 2016): Anomalies in this dataset are samples exhibiting unusual stress response elements that deviate significantly from typical cellular stress responses. These may include rare mutations or outlier stress markers.
MMP (Abdelaziz et al., 2016): Anomalous samples are identified as molecular structures with abnormal expressions of matrix metalloproteinases, often linked to extreme biological behaviors such as uncontrolled cancer metastasis.
p53 (Abdelaziz et al., 2016): The anomalies are instances with abnormal p53 protein behavior or structural deviations, which may indicate dysfunction in apoptosis or tumor suppression mechanisms.
PPAR-gamma (Abdelaziz et al., 2016): Anomalies are molecular samples exhibiting irregular activation or suppression of PPAR-gamma, which is unusual in typical metabolic and inflammatory processes.
⚫ Type II: Datasets with class-based anomaly.333 TUDataset.
This type includes six datasets (PROTEINS, ENZYMES, AIDS, DHFR, BZR, COX2) where the graphs are attributed, meaning that each node contains descriptive features. The remaining datasets are plain graphs without node-level attributes. All datasets with class-based anomalies are derived from graph classification benchmarks, and for anomaly detection tasks, the minority class is treated as anomalies. These datasets are particularly useful for evaluating models’ ability to identify rare classes or outliers within a given graph dataset.
@@ Appendix B Detailed Description of datasets in UB-GOLD
COLLAB (Yanardag & Vishwanathan, 2015): Scientific collaboration networks, representing authors and their co-authored papers in various scientific fields.
@@
IMDB-BINARY (Yanardag & Vishwanathan, 2015): Movie collaboration networks, where nodes represent actors and edges denote their co-appearance in films. This dataset helps in analyzing the collaboration patterns within the film industry.
@@
REDDIT-BINARY (Yanardag & Vishwanathan, 2015): Discussion threads from Reddit, each graph representing a thread with nodes as users and edges as interactions. It captures the structure of online discussions and their dynamics.
@@
ENZYMES (Schomburg et al., 2004): Protein tertiary structures, each graph represents an enzyme with nodes as secondary structure elements and edges as spatial adjacencies, crucial for biochemical and functional studies.
@@
PROTEINS (Dobson & Doig, 2003): Comprehensive protein structures and interaction data, where nodes represent amino acids and edges denote interactions, providing insights into protein functions and interactions.
@@
DD (Vishwanathan et al., 2010): Protein-protein interaction networks, capturing the intricate relationships between proteins within biological systems, essential for understanding cellular functions.
@@
BZR (Wu et al., 2018): Benzodiazepine receptor ligands, with nodes representing atoms and edges representing bonds, used for studying molecular interactions with benzodiazepine receptors.
@@
AIDS (Morris et al., 2020): Chemical compounds screened for antiviral activity against HIV, where each graph represents a molecule, useful in drug discovery and antiviral research.
@@
COX2 (Morris et al., 2020) : Cyclooxygenase-2 inhibitors, where nodes and edges represent molecular structures, important for studying anti-inflammatory drug properties.
@@
NCI1 (Wale et al., 2008): Chemical compounds screened for anti-cancer activity, each graph representing a molecule, used for evaluating potential anti-cancer drugs.
@@
DHFR (Morris et al., 2020): Dihydrofolate reductase inhibitors, with graphs representing molecular structures, focusing on compounds inhibiting the DHFR enzyme, vital for cancer and bacterial infection treatments.
@@
Inter-dataset shift datasets are designed by drawing ID and OOD samples from different but related datasets. This design approach allows us to simulate real-world scenarios where data distributions for both ID and OOD samples are distinct, yet share some underlying similarities.
BBBP (Martins et al., 2012): BBBP is the Blood–brain barrier penetration (BBBP) dataset includes binary labels for over 2000 compounds on their permeability properties.
BACE (Subramanian et al., 2016): BACE provides a series of human -secretases as well as their binary label, and all data are experimental values reported in the scientific literature over the last decade.
CLINTOX (Gayvert et al., 2016): The ClinTox dataset compares drugs that have been approved by the FDA with drugs that have failed in clinical trials for toxicity reasons.
LIPO (Wu et al., 2018): The full name of LIPO is Lipophilicity, which is an important feature of drug molecules that affects both membrane permeability and solubility.
FREESOLV (Mobley & Guthrie, 2014): The Free Solvation database (FreeSolv) provides experimental and calculated free energies of hydration of small molecules in water.
TOXCAST (Richard et al., 2016): ToxCast providing toxicology data for a library of compounds based on high-throughput screening and includes qualitative results of over 600 experiments on 8615 compounds.
ESOL (Delaney, 2004): ESOL is a small dataset containing water solubility data for 1128 compounds with the goal of estimating solubility from chemical structures.
MUV (Rohrer & Baumann, 2009): The Maximum Unbiased Validation (MUV) group is another benchmark dataset selected from PubChem BioAssay by applying a modified nearest neighbor analysis
TOX21 (Wu et al., 2018): The dataset contains qualitative toxicity measurements of 8014 compounds against 12 different targets, including nuclear receptors and stress response pathways.
SIDER (Altae-Tran et al., 2017): The Side Effect Resource (SIDER) is a database of marketed drugs and adverse drug reactions which measured for 1427 approved drugs.
PTC-MR (Helma et al., 2001): PTC dataset labels compounds based on their carcinogenicity where MR Indicates that their rodent is a male rat.
@@
Intra-dataset shift datasets simulate OOD conditions within a single dataset by defining shifts based on specific attributes such as size, scaffold, or assay. Unlike inter-dataset shifts, which involve drawing ID and OOD samples from different datasets, intra-dataset shifts introduce variations within the same dataset, creating challenges for models that must adapt to these changes without external data sources. Intra-dataset shifts are particularly useful for testing how well a model can handle structural or feature-based changes within a consistent data domain.
GOOD (Gui et al., 2022): A systematic benchmark specifically tailored for the graph OOD problem. We utilize two molecular datasets for OOD detection tasks: (1) GOOD-HIV is a small-scale real-world molecular dataset which aim to predict whether this molecule can inhibit HIV replication. (2) GOOD-ZINC is a real-world molecular property regression dataset from ZINC database and aims at predicting molecular solubility. Each dataset comprises two ID-OOD splitting strategies (scaffold and size), resulting in a total of 4 distinct datasets.
DrugOOD (Ji et al., 2023): This OOD benchmark is designed for AI-aided drug discovery. It includes three ID-OOD splitting strategies: assay, scaffold, and size. These strategies are applied to two measurements (IC50 and EC50), resulting in six datasets. Each dataset comprises a binary classification task aimed at predicting drug target binding affinity.
Appendix C Detailed Description of Algorithms in UB-GOLD